OpenStudy (anonymous):

help please...Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y = 1/3x - 3.

OpenStudy (anonymous):

how do you do this?

OpenStudy (anonymous):

HInt: Paralell lines have same slopes.

OpenStudy (anonymous):

You will identify the slope of the given equation first.

OpenStudy (anonymous):

hint: y=mx+b where m is the slope

OpenStudy (anonymous):

Note that y =1/3x - 3 is in the form y = mx+b, where m = slope.

OpenStudy (anonymous):

So, the slope of the given equation, will be the same on the slope of the equation we will solve.

OpenStudy (anonymous):

slope is 1/3

OpenStudy (anonymous):

After that, you will have the slope and the point, you may use the Point-Slope Form: y -y1 = (slope)(x-x1)

OpenStudy (anonymous):

that is right, slope is 1/3!

OpenStudy (anonymous):

so, plug-in 4 for y1, 6 for x1, and 1/3 for the slope.then simplify.

OpenStudy (anonymous):

example: if we have (1,2) and slope = 2/3, then: y - 2 = (2/3)(x-1).

OpenStudy (anonymous):

then, y-2 = 2/3x - 2/3

OpenStudy (anonymous):

isolating y on left side: y = 2/3x + 4/3